MINUTES OF THE PS-ISOLDE
COORDINATION MEETING held on 13
Carminati, R. Catherall, A. Dorsival, V. Fedosseev, T. Giles, J. Lettry, F.
Locci, M. Lindroos, S. Marzari, S. Meunier, T. Nilsson, F. Wenander
ISOLDE at URL: http://isolde.cern.ch
Minutes at URL: http://ps-div.web.cern.ch/ps-div/Meetings/ISOLDE/TechMtgs.html
Consolidation projects at URL: http://isolde-consolidation.web.cern.ch/isolde-consolidation/
PSMB meetings at URL: http://cern.web.cern.ch/CERN/Divisions/PS/Meetings/PSMB.html
PLC for the gas system and some measurement of potentials at the target did
not work last week. It was not essential for the run. The CO has now solved
is an intermittent fault on the scanner SC260. Please check in the shutdown.
needle scanner 670 is broken. It is very time requiring to set up the HRS in
high resolution mode without it. We urgently need it fixed.
high resolution mode of the HRS was set-up on Friday (m/deltam=2600 at 30
kV) and a substantial suppression of the contaminant Aluminum was achieved
with the slits.
and Mn run with RILIS was successful. In particular, the continuous surveillance
of ionization efficient of stable Mn in the GLM during the run with radioactive
Mn worked very well.
fiber for wavelength measurements in the infrared region is broken. It is
needed for |Sn run but a temporary solution with a broadband fiber has been
found (replacement ordered).
robot failed to position the used UC target in the shelf at the target
change last Monday. The EIC had to make an intervention to remove the target
to de-block the robot. Richard is investigating the reason for this
malfunctioning but as other shelf position could be verified to work the GPS
run was authorized to start as planned.
- The problem with the saturation of the link between the ISOLDE consoles
and the new front end computers have been resolved and the QP arrays can now
be used as before. A new front end computer has been installed with the name
PSLXGWPC and is visible in the PS ISOLDE alarm tree (for e.g. resets). This
is a LINUX PC running the standard PC front end computer software and this
solution is used elsewhere at the PC when required.
- Progress is being made on a new stepping motor system (Sven de Man) and on
the tape station software.
- Up-grade in progress on the vacuum front end computer to increase
- Sophie has at several occasions found sections at ISOLDE under atmosphere
without any particular reason. Remember to re-pump sections that you vent
even if you don't plan to continue to use or to take beam. Long exposure to
atmosphere will slow the pumping at the next occasions and waste precious
5. Radiation protection
- Take great care to sort your waste correctly in the marked containers. We
are all short of staff and we don't want to waste anyones time resorting the
bins at ISOLDE.
- The ST division is ready to implement a new solution that will resolve the
problem of target change at one separator effecting the other but it will
evidently cost some money, 5-10 kchf.
7. Off-line facilities
- A bug in the wiring of the faraday cup was found at the off-line
separator. This has some influence of the efficieny measurement made earlier
this year. Furthermore, the fact that this now is corrected means that
earlier measurements not are comparable to measurements done now.
- The first ECR prototype source is now assembled.
- You can register, submit abstracts or only pre-register
on the SAFERIB web site (http://cern.ch/saferib).
An LOI for the proposal is being written by Peter Thirolf for the september
meeting of FINUPHY.
9. NEW WEB site for minutes
- The first meeting to plan this years shutdown work will take place 4
September in the afternoon.
Lindroos from 17.08 to 25.08 (replaced by J. Lettry)
Giles from 03.09 to 13.10
20 August 2002 at 11.00
in the EP Conference Room A, building 304-1-001.